Johnson Company estimates that variable costs will be 60.00% of sales, and fixed costs will total $520,000. The selling price of the product is $4.

Compute the break-even point in (1) units and (2) dollars.

(1) Break-even sales__________ units

(2) Break even sales $__________

Assuming actual sales are $2,000,000 compute the margin of safety in (1) dollars and (2) as a ration.

(1) Margin of safety $____________

(2) Margin of safety ratio ______________%
Answer #1

Break even sales = Fixed cost/Contribution margin per unit

= 520,000/(4-60%)

= 325,000

.

Breakeven sales = 325,000 * 4

= 1,300,000

Margin of Safety = Sales - Breakeven sales

= 2,000,000 - 1,300,000

= 700,000

.

Margin of Safety ratio = 700,000/2,000,000

= 35%